40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 00:56:07 2016, gautam, Update, endtable upgrade, X endtable repopulation 
    Reply  Wed Apr 13 11:00:28 2016, Steve, Update, endtable upgrade, X endtable repopulation 
       Reply  Wed Apr 13 18:25:07 2016, gautam, Update, endtable upgrade, Lightwave health check 
          Reply  Fri Apr 15 18:38:12 2016, gautam, Update, endtable upgrade, Lightwave health check - NO IMPROVEMENT 
             Reply  Tue Apr 19 18:37:29 2016, gautam, Update, endtable upgrade, Laser swap + optical layout IR_modematch_19April2016_2.pdf
                Reply  Thu Apr 21 14:25:52 2016, gautam, Update, endtable upgrade, Green light recovered IMG_6567.JPGIMG_6568.JPGCVI_reflectivity.jpeg
                Reply  Tue Apr 26 15:22:35 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window ETMX-Tvp.jpgETMX-TvpDetail.jpg
                   Reply  Tue Apr 26 23:19:42 2016, gautam, Update, endtable upgrade, Green aligned to arm - high order mode flashes seen IMG_2219.JPGIMG_2220.JPG
                      Reply  Thu Apr 28 00:41:08 2016, gautam, Update, endtable upgrade, more progress - Transmon PD installed 
                         Reply  Fri Apr 29 00:55:46 2016, gautam, Update, endtable upgrade, green PDH locked to Xarm 
                            Reply  Mon May 2 19:14:18 2016, gautam, Update, endtable upgrade, Optical layout almost complete ETMX_wedge.pdf
                               Reply  Thu May 5 03:05:37 2016, gautam, Update, endtable upgrade, ALS status update ALS_noiseSpec_5May2016_2.pdfCoherence_5May2016.pdfimage.jpeg
                                  Reply  Thu May 5 04:05:03 2016, ericq, Update, LSC, Aux X PDH checks nom_noise.pngnom_olg.png
                                     Reply  Thu May 5 14:03:52 2016, ericq, Update, LSC, Further Aux X PDH tweaks 2016-05-05_newals.pdf
                                  Reply  Thu May 5 14:05:01 2016, rana, Update, endtable upgrade, ALS status update 
                                  Reply  Thu May 5 21:28:44 2016, gautam, Update, endtable upgrade, Innolight PZT capacitance 
                               Reply  Thu May 19 16:29:20 2016, Steve, Update, endtable upgrade, Optical layout almost complete ETMX_4x3_closed.jpgsealedETMXenclosure.jpg
                                  Reply  Fri Sep 30 19:53:07 2016, gautam, Update, endtable upgrade, X end IR pickoff fiber coupled 
                                     Reply  Wed Oct 5 16:28:10 2016, gautam, Update, endtable upgrade, EX laser power monitor PD installed 
                   Reply  Wed Apr 27 15:04:47 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window 1cETMX-Tcmp.jpg
                      Reply  Fri Apr 29 16:05:23 2016, gautam, Update, endtable upgrade, Cleaning ETMX vacuum dirty window IMG_0755.JPG
Message ID: 12073     Entry time: Wed Apr 13 00:56:07 2016     Reply to this: 12074
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: X endtable repopulation 

Over the last couple of days, I've been working on restoring the optical layout on the X-endtable. Some notes about the status as of today:

Lightwave NPRO output power

The output power from the lightwave NPRO is about 210mW (as measured with the calorimeter). This is significantly lower than the value of ~300mW reported in this elog. It may be that the laser crystal temperature has changed compared to that measurement, but the "ADJ" parameter is at 0, both today and in that measurement. The laser has also been on for more than a day now, that should be sufficient time for the crystal to equilibriate to its final operating state? Is such a large change in output power possible just because of a change in laser crystal temperature? Or did the laser really lose ~1/3rd of its output power over the last two months?

Alignment into IR Faraday, and changes to the planned layout

I've set up the layout until steering the beam through the IR faraday. The input power into the IR Faraday is ~210mW. The output power is ~186mW, after optimizing the angle of the HWP. These numbers seem consistent with what I had reported in this elog (although this was for the Innolight NPRO). The alignment looks reasonably good to the eye as well.

I've made one change to the planned layout (latest version here). Y1 is now a 2" 99% reflective for S polarization beam splitter, instead of a 1" HR mirror. I made this change because we want some light from the NPRO to be transmitted through this optic to couple into the fiber eventually, for the IR beat. I measured the transmitted power to be ~1.5mW, which is around what we were coupling into the fiber before, and should suffice now. The Lightwave NPRO datasheet (page 4) suggests that the polarization of the output of the laser is S, and the measured power before and after this optic suggests that it is working as advertised. This means that HWP 1 also has to be moved downstream (to rotate the polarization so as to maximize transmission through the IR faraday). Space constraints meant that I could not mount HWP 1 on the baseplate+3/4" OD post assembly which is what we want where possible on the new table, so for this optic, I used a 1" OD post and a fork. There may be a couple of other optics in the final layout where space constraints dictate we compromise in this way.

I've also installed beam dumps for the rejected light from the Faraday. For now, these are the old beam dumps. They looked reasonably intact. I believe we have a bunch of new beam dumps on hand as well, so these can be swapped out if deemed necessary.

Cleaning of optics

All the optics are being cleaned using first contact before being installed on the table. 

As I found out the hard way, it is not a good idea to clean small optics like half-wave plates while in their mounts. The first contact tends to bond to the frame while drying, and doesn't come off cleanly. Koji helped me clean the offending pieces (he used tweezers to manually remove the residual first contact, and then some acetone to clean up any remaining residue). Subsequently, he re-cleaned these optics, again using first contact, but this time being careful not to extend all the way out to the edge of the optic. The idea is to cover as much area as possible with first contact, while staying clear of the edge. This approach worked reasonably well.


The next major step is to achieve optimal alignment into the doubler. I've placed the doubler on the table in it's approximate final position, I wanted to make sure the enclosure support wasn't in the way (it isn't). The cable from the oven won't run all the way to the Thorlabs temperature controller in it's usual place, we need to either extend the cable, or figure out a new place where we can keep the temperature controller.

ELOG V3.1.3-